# The effect of accretion on scalar superradiant instability

### v1

@@ IV.2 Evolution of the BH and dominate modes
The analysis is quite similar to that from Eq. (45) to Eq. (54). Between $t_{1}^{(0ll)}$ and $t_{1}^{(0,l+1,l+1)}$, the BH evolves along the $\{0,l,l\}$ Regge curve. For most of this time range, the evolution is dominated by the accretion and the $\{0,l,l\}$ mode in Eq. (66), with other terms safely neglected. On the other hand, the $\dot{a}_{*}$ can also be calculated using the Eq. (10). Combining these two expressions of $\dot{a}_{*}$, one could solve $\dot{E}_{\text{SR}}^{(0ll)}$ as a function of $\alpha$. Then taking GW emission in Eq. (11c) as a perturbation and assuming $\alpha$ increases with time as in Eq. (19), one could further solve $M_{\text{s,Regge}}^{(0ll)}$ and the mass of the $\{0,l,l\}$ mode in this time range
$\begin{split}M_{\text{s}}^{(0ll)}(t)&=M_{\text{s}}^{(0ll)}(t_{1}^{(0ll)})+M_{\text{s,Regge}}^{(0ll)}(\alpha)\\ &\hskip 14.22636pt-M_{\text{s,Regge}}^{(0ll)}(\alpha_{1}^{(0ll)})+\delta M_{\text{s}}^{(0ll)}(t).\end{split}$ (71)
The expressions of $M_{\text{s,Regge}}^{(0ll)}$ and $\delta M_{\text{s}}$ for $l\leq 7$ are given in Eqs. (101) and (102), respectively.
The value of $\alpha_{2}^{(0ll)}$ can be determined with $d\left(M_{\text{s}}^{(0ll)}/M\right)/dt=0$ at $t_{2}^{(0ll)}$, which gives an equation like Eq. (50), with the superscript $(011)$ replaced by $(0ll)$. Then one could solve $\alpha_{2}^{(0ll)}$ from this equation. Using the value of $\alpha_{2}^{(0ll)}$, one could estimate $t_{2}^{(0ll)}$ by assuming the BH mass grows as in Eq. (19) from $t=0$ to $t_{2}^{(0ll)}$. A more accurate estimate can be obtained by inserting the $\dot{E}_{\text{Regge}}^{(0ll)}$ into Eq. (11a) and solving the differential equation, in the same manner as Eqs. (52) and (53)
$M\approx\frac{2l\alpha_{1}^{(0ll)}e^{(t-t_{1}^{(0ll)})/\tau}/\mu}{2l+3\sqrt{6}\alpha_{1}^{(0ll)}\left[e^{(t-t_{1}^{(0ll)})/\tau}-1\right]}.$ (72)
Finally, one can solve out $t_{2}^{(0ll)}$ with this expression of $M(t)$ and the value of $\alpha_{2}^{(0ll)}$. The obtained $\alpha_{2}^{(0ll)}$ are listed in Eqs. (104) for $l\leq 7$. Inserting these values into Eq. (71), the maximum mass ratio of the $\{0,l,l\}$ mode could be calculated.
@@
IV.2.2 $a_{*0}=0$
In the case of $a_{*0}<a_{*\text{c}}^{(0,l_{\text{c}}l_{\text{c}})}$, one first identify the value of $l^{\prime}$ with Eq. (65), which is $\infty$ for $a_{*0}=0$.
To get $l_{\text{acc}}$, we calculate $\alpha_{0^{\prime}}^{(0ll)}$ with $l$ increasing from $l_{\text{c}}$ to $l^{\prime}$. Here $\alpha_{0^{\prime}}^{(0ll)}$ is the mass coupling at the time when the BH spin reaches the local maximum before transiting to the $\{0,l,l\}$ Regge curve if only accretion and the $\{0,l,l\}$ mode are present. For $l<l^{\prime}$, the calculation of $\alpha_{0^{\prime}}^{(011)}$ starts from the intersection of the accretion trajectory and the $\{0,l,l\}$ Regge curve, which is given in Eq. (59). We start from $l=m=1$, the mass coupling at the intersection is $\alpha_{\text{int}}^{(011)}=0.116$. The mass of the cloud can be estimated using Eq. (37). For our example, one gets $M_{\text{s,int}}^{(011)}=6\times 10^{-25987594}\approx 0$. Then taking the intersection as the new initial time, one could eventually apply Eq. (58) and get $\alpha_{0^{\prime}}^{(011)}=0.117$.
Similarlly, one could calculate the $M_{\text{s}}^{(0ll)}$ and $\alpha_{0^{\prime}}^{(0ll)}$ for any $\{0,l,l\}$ modes. Their expressions for $l\leq 7$ are listed in Eqs. (88) and (90). For our example, we have $\alpha_{0^{\prime}}^{(022)}=0.107$ and $\alpha_{0^{\prime}}^{(033)}=0.232$. Therefore, we conclude $l_{\text{acc}}=2$.
Then one could choose the intersection of the accretion trajectory and the $\{0,2,2\}$ Regge curve as $t=0$, all the steps explained in Sec. IV.2.1 can be applied for the later evolution. The comparison of the analytical approximation with the numerical results is shown in Fig. 6(k).
@@ IV.2.2 $a_{*0}=0$
Table 2: Comparison of the analytical estimates to the numerical results obtained from solving (11) for essential quantities. Initial parameters are identical to those used for the solid lines in Fig. 7.
IV.3 Subdominant modes
The subdominant modes with indices $\{1,l,l\}$ are important for GW beat signature. In this part, we explain the calculation of the maximum mass and the lifetime of these modes. The critical BH spin of the subdominant mode $a_{*}^{(1ll)}$ can be directly calculated with Eq. (10). It is a little larger than that of the dominant mode with the same value of $l$. As a result, the $\{1,l,l\}$ Regge curve is a bit above the $\{0,l,l\}$ one. Then at the beginning of the $l$-stage when the BH spin quickly decreases, the evolution first reaches the $\{1,l,l\}$ curve. It could follow along the $\{1,l,l\}$ curve for a while before stepping down to the $\{0,l,l\}$ curve. It is also likely that the evolution quickly crosses the $\{1,l,l\}$ curve and merges to the $\{0,l,l\}$ curve. In both cases, the growth time of the subdominant mode is shorter than the dominant one. In this work, we assume the initial mass of the subdominant mode is not unnaturally large. Then one could conclude that the mass of the subdominant mode is always much smaller than the corresponding dominant mode for $l\leq 2$ since the superradiance rate of the former is always smaller than the latter. For $l\geq 3$, there exists a “overtone mixing” phenomenon where the superradiance rate of the subdominant mode is comparable to, or even larger than, the dominant mode Siemonsen:2019ebd . Interestingly, the above analytical approximations still work for the correct order of magnitude. We will discuss this scenario in Sec.V. In the rest of this section, we assume the subdominant mode always has a smaller mass than the corresponding dominant mode.
@@ IV.3 Subdominant modes
We first study the maximum mass of the subdominant mode. To be consistent in the notation with the derivation of the dominant mode, we define this time as $t_{2}^{(111)}$. Since it must be smaller than $t_{2}^{(011)}$, the GW emission can be safely ignored before $t_{2}^{(111)}$. We first assume it is larger than $t_{1}^{(011)}$ when the BH spin reaches the local minimum. If both $\{0,1,1\}$ and $\{1,1,1\}$ modes are included, the left side of Eq. (45) should be changed to $\dot{E}_{\text{SR}}^{(011)}+\dot{E}_{\text{SR}}^{(111)}$. At time $t_{2}^{(111)}$, the BH spin equals to $a_{*\text{c}}^{(111)}$ given in Eq. (10), causing $\dot{E}_{\text{SR}}^{(111)}=0$. On the other hand, $\dot{E}_{\text{SR}}^{(011)}=2M_{\text{s}}^{(011)}\Gamma_{011}$, in which $\Gamma_{011}$ is given in Eq. (25) and $M_{\text{s}}^{(011)}$ can be estimated by
$\begin{split}M_{\text{s}}^{(011)}(\alpha_{2}^{(111)})\approx&{M}_{\mathrm{s,1}}^{(011)}+{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2}^{(111)})\\ &-{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{1}^{(011)}),\end{split}$ (74)
with ${M}_{\mathrm{s,Regge}}^{(011)}$ given in Eq. (46). Then one could solve for the value of $\alpha_{2}^{(111)}$ at the LO of $\alpha_{0}$
$\alpha^{(111)}_{2}$ $\approx\left(\frac{864}{5\mu{\tau}_{\mathrm{acc}}}\right)^{1/11}.$ (75)
@@
${M}_{\mathrm{s,max}}^{(1ll)}$ $\approx{M}_{\mathrm{s,0}}^{(1ll)}\left[\frac{{M}_{\mathrm{s}}^{(0ll)}(t_{\text{2}}^{(1ll)})}{{M}_{\mathrm{s,0}}^{(0ll)}}\right]^{\beta_{1l}/\beta_{0l}},$ (76)
@@
$\beta_{nl}\equiv\frac{(n+2l+1)!}{n!(n+l+1)^{2l+4}},$ (77)
@@
After reaching the maximum mass, the subdominant mode enters the decaying phase and starts to shrink in size. We define $t_{3}^{(1ll)}$ as the time when the $\{1,l,l\}$ mode mass decreases to $1/e$ of its peak value. Two scenarios exist for estimating $t_{3}^{(1ll)}$. If the evolution does not follow the $\{0,l,l\}$ Regge curve, the accretion can be ignored. Then $t_{3}^{(1ll)}-t_{1}^{(0ll)}$ could be calculated in the same way as $\tau_{\text{life}}^{(111)}$ explained in Sec. II.2.2. The results with $l\leq 7$ are listed in Eqs. (86). On the other hand, if the evolution follows the $\{0,l,l\}$ Regge curve, the growth of the BH mass is important. One starts from Eq. (11c) with GW emission ignored. Using $a_{*}=a_{*\text{c}}^{(0ll)}$ given in Eq. (10) and the exponential form of $M$ given in Eq. (19), one could then integrate Eq. (11c) from $t_{1}^{(0ll)}$ to $t_{3}^{(1ll)}$ and obtain an expression similar to Eq. (39). Then it is straightforward to calculate the mass coupling $\alpha_{3}^{(111)}$ at time $t_{3}^{(111)}$. Finally, one could solve for $t_{3}^{(111)}$ from the evolution of $M$ on the $\{0,l,l\}$ Regge curve. The calculation of the $M(t)$ on any $\{0,l,l\}$ Regge curve is explained in the previous subsection. For $l=1$, one arrives at Eq. (54) with $\alpha_{2}$ replaced by $\alpha_{3}^{(011)}$.
IV.4 GW emission
@@
We finally discuss the GW emission of each $l$-stage and the plausibility of observation by future GW interferometers. In the example presented in Fig. 7, one could observe that the GW emission flux in the $l=2$ and $l=3$ stages are comparable to that in the $l=1$ stage with accretion. It is very different from the case without accretion, in which the GW flux of the $l>1$ stage is many decades smaller Guo:2022mpr .
@@
To study how accretion affects the GW emission with other parameters, the GW emission fluxes of the dominant modes with $l\leq 3$ and their timescales are plotted as functions of $M_{0}\mu$ in Fig. 8 (a) and (b), with four different accretion rates ${f}_{\mathrm{Edd}}$. The analytical approximations of the BH mass $M$, cloud mass $M_{\text{s}}^{(0ll)}(t)$, and different time scales are utilized. We set the scalar mass as $\mu=10^{-17}$ eV, the initial BH spin as $a_{*}=0.998$, and the initial condensate masses for all dominant and subdominant modes as $10^{-5}M_{0}$. In panel (a), the horizontal line denotes the flux at which the characteristic strain equals the noise characteristic strain of the LISA, with the source located at redshift $z=0.1$, and the observation time as 4 years. The calculations are explained in Refs. Brito:2017zvb ; Guo:2022mpr .
@@
From Fig. 8 (a) and (b), one observes that accretion greatly enhances the GW emission flux, and at the same time reduces the GW duration time for small $M_{0}\mu$. With the BH mass $M$ fixed, the small-mass scalars which are undetectable via superradiance without accretion can be observed by LISA with $f_{\text{Edd}}$ as small as 0.01. This finding indicates future GW interferometers like LISA can search for a much larger range of the scalar mass than previously thought. Moreover, it is generally recognized in the literature that each BH undergoes boson annihilation only once in its cosmic history, as subsequent annihilation signals are significantly weaker Brito:2017zvb ; Brito:2017wnc . This is not generally true with accretion. On the one hand, a BH could have multiple boson annihilation in its cosmic history. For example, the lifetimes of the GW emission with $f_{\text{Edd}}=1$ and $\alpha_{0}=0.25$ are $2.18\times 10^{5}\;\mathrm{yr}$, $4.94\times 10^{7}\;\mathrm{yr}$ and $3.27\times 10^{7}\;\mathrm{yr}$ for $l=1,2,3$, respectively. The numbers are $2.18\times 10^{5}\;\mathrm{yr}$, $4.89\times 10^{8}\;\mathrm{yr}$, and $3.41\times 10^{8}\;\mathrm{yr}$ if $f_{\text{Edd}}$ is changed to 0.1. On the other hand, the GW emission flux with $l\geq 2$ modes could also be observed. This observation could be important in the study of stochastic GW background. We leave such phenomenological studies in future publications.
@@
In Fig. 8 (c) and (d), we further study the GW beat signal strength and its duration. The analytical approximations for the former are listed in App. A, with more details on GW beats given in Ref. Guo:2022mpr . The horizontal line in panel (c) is the same as that in panel (a). Same as the GW flux, the accretion also enhances the beat signal and reduces its duration in each $l$-stage. The beat signal without accretion has been carefully studied in Ref. Guo:2022mpr . With accretion, the enhanced beat signal could greatly facilitate the search for light scalars.
@@
V Summary and Discussion
In this work, we provided a thorough analysis of how accretion affects the evolution of a BH-condensate system with scalar superradiance and GW emission. Despite the coexistence of multiple modes with different values of $n$ and $l$, the evolution has a simple and universal pattern. With the small mass-coupling $\alpha=M\mu$, we obtained analytical approximations of all the important quantities. Compared with numerical results, these approximations agree reasonably well (see Tables 1 and 2). We found that accretion could greatly speed up the mass accumulation and the GW emission, resulting in stronger and shorter GW signal compared to the case without accretion.

### v2

@@ 4.2 Evolution of the BH and dominate modes
The analysis is quite similar to that from Eq. (3.12) to Eq. (3.21). Between $t_{1}^{(0ll)}$ and $t_{1}^{(0,l+1,l+1)}$, the BH evolves along the $\{0,l,l\}$ Regge curve. For most of this time range, the evolution is dominated by the accretion and the $\{0,l,l\}$ mode in Eq. (4.6), with other terms safely neglected. On the other hand, the $\dot{a}_{*}$ can also be calculated using the Eq. (2.10). Combining these two expressions of $\dot{a}_{*}$, one could solve $\dot{E}_{\text{SR}}^{(0ll)}$ as a function of $\alpha$. Then taking GW emission in Eq. (2.11c) as a perturbation and assuming $\alpha$ increases with time as in Eq. (2.19), one could further solve $M_{\text{s,Regge}}^{(0ll)}$ and the mass of the $\{0,l,l\}$ mode in this time range
$M_{\text{s}}^{(0ll)}(t)=M_{\text{s}}^{(0ll)}(t_{1}^{(0ll)})+M_{\text{s,Regge}}^{(0ll)}(\alpha)-M_{\text{s,Regge}}^{(0ll)}(\alpha_{1}^{(0ll)})+\delta M_{\text{s}}^{(0ll)}(t).$ (4.11)
The expressions of $M_{\text{s,Regge}}^{(0ll)}$ and $\delta M_{\text{s}}$ for $l\leq 7$ are given in Eqs. (A.18) and (A.19), respectively.
The value of $\alpha_{2}^{(0ll)}$ can be determined with $d\left(M_{\text{s}}^{(0ll)}/M\right)/dt=0$ at $t_{2}^{(0ll)}$, which gives an equation like Eq. (3.17), with the superscript $(011)$ replaced by $(0ll)$. Then one could solve $\alpha_{2}^{(0ll)}$ from this equation. Using the value of $\alpha_{2}^{(0ll)}$, one could estimate $t_{2}^{(0ll)}$ by assuming the BH mass grows as in Eq. (2.19) from $t=0$ to $t_{2}^{(0ll)}$. A more accurate estimate can be obtained by inserting the $\dot{E}_{\text{Regge}}^{(0ll)}$ into Eq. (2.11a) and solving the differential equation, in the same manner as Eqs. (3.19) and (3.20)
$M\approx\frac{2l\alpha_{1}^{(0ll)}e^{(t-t_{1}^{(0ll)})/\tau}/\mu}{2l+3\sqrt{6}\alpha_{1}^{(0ll)}\left[e^{(t-t_{1}^{(0ll)})/\tau}-1\right]}.$ (4.12)
Finally, one can solve out $t_{2}^{(0ll)}$ with this expression of $M(t)$ and the value of $\alpha_{2}^{(0ll)}$. The obtained $\alpha_{2}^{(0ll)}$ are listed in Eqs. (A.21) for $l\leq 7$. Inserting these values into Eq. (4.11), the maximum mass ratio of the $\{0,l,l\}$ mode could be calculated.
@@
4.2.2 $a_{*0}=0$
In the case of $a_{*0}<a_{*\text{c}}^{(0,l_{\text{c}}l_{\text{c}})}$, one first identify the value of $l^{\prime}$ with Eq. (4.5), which is $\infty$ for $a_{*0}=0$.
To get $l_{\text{acc}}$, we calculate $\alpha_{0^{\prime}}^{(0ll)}$ with $l$ increasing from $l_{\text{c}}$ to $l^{\prime}$. Here $\alpha_{0^{\prime}}^{(0ll)}$ is the mass coupling at the time when the BH spin reaches the local maximum before transiting to the $\{0,l,l\}$ Regge curve if only accretion and the $\{0,l,l\}$ mode are present. For $l<l^{\prime}$, the calculation of $\alpha_{0^{\prime}}^{(011)}$ starts from the intersection of the accretion trajectory and the $\{0,l,l\}$ Regge curve, which is given in Eq. (3.26). We start from $l=m=1$, the mass coupling at the intersection is $\alpha_{\text{int}}^{(011)}=0.116$. The mass of the cloud can be estimated using Eq. (3.4). For our example, one gets $M_{\text{s,int}}^{(011)}=6\times 10^{-25987594}\approx 0$. Then taking the intersection as the new initial time, one could eventually apply Eq. (3.25) and get $\alpha_{0^{\prime}}^{(011)}=0.117$.
Similarlly, one could calculate the $M_{\text{s}}^{(0ll)}$ and $\alpha_{0^{\prime}}^{(0ll)}$ for any $\{0,l,l\}$ modes. Their expressions for $l\leq 7$ are listed in Eqs. (A.11) and (A.13). For our example, we have $\alpha_{0^{\prime}}^{(022)}=0.107$ and $\alpha_{0^{\prime}}^{(033)}=0.232$. Therefore, we conclude $l_{\text{acc}}=2$.
Then one could choose the intersection of the accretion trajectory and the $\{0,2,2\}$ Regge curve as $t=0$, all the steps explained in Sec. 4.2.1 can be applied for the later evolution. The comparison of the analytical approximation with the numerical results is shown in Fig. 6(k).
@@ 4.2.2 $a_{*0}=0$
Table 2: Comparison of the analytical estimates to the numerical results obtained from solving (2.11) for essential quantities. Initial parameters are identical to those used for the solid lines in Fig. 7.
4.3 Subdominant modes
The subdominant modes with indices $\{1,l,l\}$ are important for GW beat signature. In this part, we explain the calculation of the maximum mass and the lifetime of these modes. The critical BH spin of the subdominant mode $a_{*}^{(1ll)}$ can be directly calculated with Eq. (2.10). It is a little larger than that of the dominant mode with the same value of $l$. As a result, the $\{1,l,l\}$ Regge curve is a bit above the $\{0,l,l\}$ one. Then at the beginning of the $l$-stage when the BH spin quickly decreases, the evolution first reaches the $\{1,l,l\}$ curve. It could follow along the $\{1,l,l\}$ curve for a while before stepping down to the $\{0,l,l\}$ curve. It is also likely that the evolution quickly crosses the $\{1,l,l\}$ curve and merges to the $\{0,l,l\}$ curve. In both cases, the growth time of the subdominant mode is shorter than the dominant one. In this work, we assume the initial mass of the subdominant mode is not unnaturally large. Then one could conclude that the mass of the subdominant mode is always much smaller than the corresponding dominant mode for $l\leq 2$ since the superradiance rate of the former is always smaller than the latter. For $l\geq 3$, there exists a “overtone mixing” phenomenon where the superradiance rate of the subdominant mode is comparable to, or even larger than, the dominant mode [33]. Interestingly, the above analytical approximations still work for the correct order of magnitude. We will discuss this scenario in Sec.5. In the rest of this section, we assume the subdominant mode always has a smaller mass than the corresponding dominant mode.
@@ 4.3 Subdominant modes
We first study the maximum mass of the subdominant mode. To be consistent in the notation with the derivation of the dominant mode, we define this time as $t_{2}^{(111)}$. Since it must be smaller than $t_{2}^{(011)}$, the GW emission can be safely ignored before $t_{2}^{(111)}$. We first assume it is larger than $t_{1}^{(011)}$ when the BH spin reaches the local minimum. If both $\{0,1,1\}$ and $\{1,1,1\}$ modes are included, the left side of Eq. (3.12) should be changed to $\dot{E}_{\text{SR}}^{(011)}+\dot{E}_{\text{SR}}^{(111)}$. At time $t_{2}^{(111)}$, the BH spin equals to $a_{*\text{c}}^{(111)}$ given in Eq. (2.10), causing $\dot{E}_{\text{SR}}^{(111)}=0$. On the other hand, $\dot{E}_{\text{SR}}^{(011)}=2M_{\text{s}}^{(011)}\Gamma_{011}$, in which $\Gamma_{011}$ is given in Eq. (2.25) and $M_{\text{s}}^{(011)}$ can be estimated by
$M_{\text{s}}^{(011)}(\alpha_{2}^{(111)})\approx$ ${M}_{\mathrm{s,1}}^{(011)}+{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{2}^{(111)})-{M}_{\mathrm{s,Regge}}^{(011)}(\alpha_{1}^{(011)}),$ (4.14)
with ${M}_{\mathrm{s,Regge}}^{(011)}$ given in Eq. (3.13). Then one could solve for the value of $\alpha_{2}^{(111)}$ at the LO of $\alpha_{0}$
$\alpha^{(111)}_{2}$ $\approx\left(\frac{864}{5\mu{\tau}_{\mathrm{acc}}}\right)^{1/11}.$ (4.15)
@@
${M}_{\mathrm{s,max}}^{(1ll)}$ $\approx{M}_{\mathrm{s,0}}^{(1ll)}\left[\frac{{M}_{\mathrm{s}}^{(0ll)}(t_{\text{2}}^{(1ll)})}{{M}_{\mathrm{s,0}}^{(0ll)}}\right]^{\beta_{1l}/\beta_{0l}},$ (4.16)
@@
$\beta_{nl}\equiv\frac{(n+2l+1)!}{n!(n+l+1)^{2l+4}},$ (4.17)
@@
After reaching the maximum mass, the subdominant mode enters the decaying phase and starts to shrink in size. We define $t_{3}^{(1ll)}$ as the time when the $\{1,l,l\}$ mode mass decreases to $1/e$ of its peak value. Two scenarios exist for estimating $t_{3}^{(1ll)}$. If the evolution does not follow the $\{0,l,l\}$ Regge curve, the accretion can be ignored. Then $t_{3}^{(1ll)}-t_{1}^{(0ll)}$ could be calculated in the same way as $\tau_{\text{life}}^{(111)}$ explained in Sec. 2.2.2. The results with $l\leq 7$ are listed in Eqs. (A.9). On the other hand, if the evolution follows the $\{0,l,l\}$ Regge curve, the growth of the BH mass is important. One starts from Eq. (2.11c) with GW emission ignored. Using $a_{*}=a_{*\text{c}}^{(0ll)}$ given in Eq. (2.10) and the exponential form of $M$ given in Eq. (2.19), one could then integrate Eq. (2.11c) from $t_{1}^{(0ll)}$ to $t_{3}^{(1ll)}$ and obtain an expression similar to Eq. (3.6). Then it is straightforward to calculate the mass coupling $\alpha_{3}^{(111)}$ at time $t_{3}^{(111)}$. Finally, one could solve for $t_{3}^{(111)}$ from the evolution of $M$ on the $\{0,l,l\}$ Regge curve. The calculation of the $M(t)$ on any $\{0,l,l\}$ Regge curve is explained in the previous subsection. For $l=1$, one arrives at Eq. (3.21) with $\alpha_{2}$ replaced by $\alpha_{3}^{(011)}$.
4.4 GW emission
@@
We finally discuss the GW emission of each $l$-stage and the plausibility of observation by future GW interferometers. In the example presented in Fig. 7, one could observe that the GW emission flux in the $l=2$ and $l=3$ stages are comparable to that in the $l=1$ stage with accretion. It is very different from the case without accretion, in which the GW flux of the $l>1$ stage is many decades smaller [16].
@@
To study how accretion affects the GW emission with other parameters, the GW emission fluxes of the dominant modes with $l\leq 3$ and their timescales are plotted as functions of $M_{0}\mu$ in Fig. 8 (a) and (b), with four different accretion rates ${f}_{\mathrm{Edd}}$. The analytical approximations of the BH mass $M$, cloud mass $M_{\text{s}}^{(0ll)}(t)$, and different time scales are utilized. We set the scalar mass as $\mu=10^{-17}$ eV, the initial BH spin as $a_{*}=0.998$, and the initial condensate masses for all dominant and subdominant modes as $10^{-5}M_{0}$. In panel (a), the horizontal line denotes the flux at which the characteristic strain equals the noise characteristic strain of the LISA, with the source located at redshift $z=0.1$, and the observation time as 4 years. The calculations are explained in Refs. [14, 16].
@@
From Fig. 8 (a) and (b), one observes that accretion greatly enhances the GW emission flux, and at the same time reduces the GW duration time for small $M_{0}\mu$. With the BH mass $M$ fixed, the small-mass scalars which are undetectable via superradiance without accretion can be observed by LISA with $f_{\text{Edd}}$ as small as 0.01. This finding indicates future GW interferometers like LISA can search for a much larger range of the scalar mass than previously thought. Moreover, it is generally recognized in the literature that each BH undergoes boson annihilation only once in its cosmic history, as subsequent annihilation signals are significantly weaker [14, 13]. This is not generally true with accretion. On the one hand, a BH could have multiple boson annihilation in its cosmic history. For example, the lifetimes of the GW emission with $f_{\text{Edd}}=1$ and $\alpha_{0}=0.25$ are $2.18\times 10^{5}\;\mathrm{yr}$, $4.94\times 10^{7}\;\mathrm{yr}$ and $3.27\times 10^{7}\;\mathrm{yr}$ for $l=1,2,3$, respectively. The numbers are $2.18\times 10^{5}\;\mathrm{yr}$, $4.89\times 10^{8}\;\mathrm{yr}$, and $3.41\times 10^{8}\;\mathrm{yr}$ if $f_{\text{Edd}}$ is changed to 0.1. On the other hand, the GW emission flux with $l\geq 2$ modes could also be observed. This observation could be important in the study of stochastic GW background. We leave such phenomenological studies in future publications.
@@
In Fig. 8 (c) and (d), we further study the GW beat signal strength and its duration. The analytical approximations for the former are listed in App. A, with more details on GW beats given in Ref. [16]. The horizontal line in panel (c) is the same as that in panel (a). Same as the GW flux, the accretion also enhances the beat signal and reduces its duration in each $l$-stage. The beat signal without accretion has been carefully studied in Ref. [16]. With accretion, the enhanced beat signal could greatly facilitate the search for light scalars.
@@
5 Summary and Discussion
In this work, we provided a thorough analysis of how accretion affects the evolution of a BH-condensate system with scalar superradiance and GW emission. Despite the coexistence of multiple modes with different values of $n$ and $l$, the evolution has a simple and universal pattern. Our main findings are summarized below. In theoretical aspect:
  • We considered, for the first time, the capture of photons from the accretion disk by the BH during superradiant evolution. Despite the presence of superradiance, the BH spin still does not exceed its limit value $a_{*\text{lim}}\simeq 0.998$.
  • With the small mass-coupling $\alpha=M\mu$, we obtained analytical approximations of all the important quantities, incorporating the effects of superradiance, accretion and GW emission. These formulas can be readily applied in relevant phenomenological studies.
  • This work serves as a theoretical baseline for more complex scenarios, such as including non-gravitational interactions between the scalar field and Standard Model particles.
In phenomenological aspect, we found:
  • Accretion could greatly speed up the mass accumulation and the GW emission, resulting in stronger and shorter GW signal compared to the case without accretion.
  • GWs emitted by modes with $l>1$ should also be taken into account when studying the stochastic GW background and conducting targeted GW searches, even though they are often neglected in the literature.